# A voice for minorities: diversity in approval-based committee elections under incomplete or inaccurate information

### v1

@@ 4.1 Upper bound for the unconstrained setting
ALGORITHM 1 Greedy
Lemma 4.1.
Algorithm 1 is $(1-\frac{1}{\operatorname{\text{e}}})$-approximate for the Max Cover, and runs in time $\mathcal{O}(kmn)$.
This is a well known result, see e.g. [16]. We provide our own proof of Lemma 4.1 in the appendix. Writing $\operatorname{\textsc{gr}}(k)$ for the Max Cover score of the greedy algorithm after termination, we will assume that $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\geq k$; there exist at least $k$ voters that approve a candidate. We call such instances, for which $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\geq k$ is true, non-trivial instances. After running the algorithm, thereby gathering the query responses, we can verify efficiently whether an instance is non-trivial.
Proposition 4.2.
On any non-trivial instance, $\operatorname{\textsc{gr}}(k)\geq k$.
For the incomplete information setting, that we discuss in detail in the next section, we estimate approval information of the total population, based on the information of only few voters. This can lead to errors. Write $\operatorname{\textsc{gr}}\text{-}\epsilon$ for an adaptation of Algorithm 2 that, in each iteration, might choose any candidate that yields the largest increase in score up to $-\epsilon$. We can compute the approximation ratio in a similar fashion as before in Lemma 4.1.
Lemma 4.3.
For any non-trivial instance, greedy-$\epsilon$ is $\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)\cdot\left(\frac{1}{1+\epsilon(1-\frac{1}{\operatorname{\text{e}}})}\right)\right)$-approximate.
The greedy query algorithm
We now turn to the incomplete information setting. Recall that here we assume there is an underlying unknown distribution $d$ of voter preferences and we write $A:V\to\mathcal{P}(C)$ for the random variable that maps a voter $i$ to its approval set $A(i)$, and is distributed according to $d$. In the query model, we repeatedly sample sets (of size $\ell\leq n$) of voters independently from $d$ and ask these voters for their votes on some query set $Q\subseteq C$ (of size $t$). Querying any set $Q\subseteq C$, on all voters, the responses allow us to compute, for any $S\subseteq Q$,
$p_{S}:=\mathbb{P}(A\cap S\neq\emptyset)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(A(i)\cap S\neq\emptyset);$ (2)
the Chamberlin-Courant score of $S$. We can now write $\Delta(W,c^{\prime}):=p_{W\cup\{c^{\prime}\}}-p_{W}$, and we can compute this with complete information on a query containing $W\cup\{c^{\prime}\}$. Sampling $\ell$ voters from $d$ independently and presenting them with query set $Q\subseteq C$, we can compute, for every subset $S\subseteq Q$,
$\hat{p}_{S}:=\frac{1}{\ell}\sum_{i=1}^{\ell}\mathds{1}(\{A(i)\cap S\neq\emptyset\}),$ (3)
that estimates $p_{S}$. We write $\hat{\Delta}(W,c^{\prime})=\hat{p}_{W\cup\{c^{\prime}\}}-\hat{p}_{W}$ for an estimate of $\Delta(W,c^{\prime})$ based on values $\hat{p}$. Using this querying approach, we obtain Algorithm 2. Theorem 4.4 gives an upper bound on the amount of query sets required to run Algorithm 2.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C\subseteq\mathcal{P}(V)$ of $m$ candidates, query size $t$, with $m\geq t>k$, and $\gamma\in(0,1),\ \delta\in(0,1]$.
Output: Committee $W\subseteq C$ of size $k$.
Let $W=[\ ]$ be an empty list;
for $i=1,\ldots,k$, do
  1. 1.
    Construct a set of query sets $\mathcal{Q}=\{Q_{i}\}_{i}$ with $Q_{i}\subseteq C$ and $|Q_{i}|=t$ for all $i$, and such that $W\subseteq\bigcap\mathcal{Q}$ and $\bigcup\mathcal{Q}=C$, and present each query set to $\ell$ voters.
  2. 2.
    For all $c^{\prime}\notin W$, determine $\hat{\Delta}(W,c^{\prime})$ as an estimate of $\Delta(W,c^{\prime})$ using $\ell$ voters and query set $Q$ containing $W\cup c^{\prime}$.
  3. 3.
    Append $W$ with $\operatorname*{argmax}_{c^{\prime}\notin W}\,\hat{\Delta}(W,c^{\prime})$.
      end for
ALGORITHM 2 Greedy-queries
Theorem 4.4.
For any non-trivial instance, any $\delta>0,\,\gamma\in(0,1)$, and $k<t\leq m$, with probability at least $1-\delta$, Algorithm 2 is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover using
$k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\in\mathcal{O}\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)$
queries.
Example 4.5.
For $m=1000,\,\gamma=0.85,\,\delta=0.05,\,k=8$ and $t=20$, the upper bound of Theorem 4.4 gives 432 920 queries.
4.2 Upper bound for a setting with matroid constraints
@@
We can imagine situations where we want to place restrictions on the set of possible committees to be elected. For instance, perhaps we want to include ‘external’ diversity requirements, such as ‘the committee must contain at least $x$ candidates of category $Y$’. In practice, $x$ would be a natural number and $Y$ would be a category like gender, race, etc. In online deliberation settings, these attributes would extend to users. Or perhaps we want to place a restriction on how many comments can be selected that are submitted by one and the same person. For such purposes, we make a selection of which committees are “allowed” to be elected, and collect them in a set $\mathcal{M}$. This is fundamentally different from what we have so far meant with diversity, which was just defined in terms of the approval profile. It is not obvious that any algorithm adapts well to this restriction. Fortunately, [12] prove that, when $\mathcal{M}$ is a matroid, we can mend this problem.
Definition 4.6 (matroid).
@@
This last property guarantees that all maximal independent sets have the same cardinality. We call such sets bases, and their common cardinality is called the rank of the matroid.
Example 4.7 (uniform matroid).
Consider =$\mathcal{M}=\{W\subset C:|W|\leq k\}$. This is the uniform matroid of rank $k$: 1) $|\emptyset|=0\leq k$, 2) for any set $W\in\mathcal{M}$ with $|W|=j\leq k$, for any subset $W^{\prime}\subseteq W$, $|W^{\prime}|\leq|W|\leq k$, and 3) for any two feasible sets $W,\ W^{\prime}\in\mathcal{M}$ with $|W^{\prime}|<|W|$, we will have, for any $c\in W\setminus W^{\prime}$, that $|W^{\prime}\cup\{c\}|\leq k$.
Thus, taking for $\mathcal{M}$ the uniform matroid of rank $k$, we retrieve the original problem. This shows the problem with matroid constraints is a generalization of the original problem, hence more difficult.
Example 4.8 (a matroid implementing quota).
Define a set of bases $M=\{W\subseteq C:|W|=k,\ \ |W\cap Y|=\frac{1}{2}k,\ |W\cap X|=\frac{1}{2}k\}$ and consider $\mathcal{M}=M\bigcup\ \{V:V\subseteq W,\ W\in M\}$. This is a matroid of rank $k$.
In Example 4.8, only the bases elements comply with the quota. This is not a problem, we show next.
Proposition 4.9.
For any instance $(V,C,A,k)$ and any matroid $\mathcal{M}$ of rank $k$, no cover of size $\ell<k$ will attain a unique maximal Max Cover score.
Thus, there is always basis among those covers with the highest score. When an algorithm does output a cover $L$ of size $\ell<k$, we can take any basis $K$ and add elements from $K$ to $L$ until it has rank $k$.
@@
As done above, we start with the complete information setting. Here, a classical (or, oblivious) local search algorithm (for Max Cover) typically starts with an arbitrary (tentative) solution and, in each iteration, swaps one or multiple of the tentatively elected sets for tentatively unelected sets in order to improve the objective function value. It thus searches in the local neighborhood of the solution. It stops when no local improvement is possible anymore, which could in principle take exponentially many steps. As described in [12], this running time problem can be circumvented. We can use approximate local search, which allows us to bound the running time at a small cost in the approximation ratio, see also [27]. In practice, this means we stop the algorithm sooner, namely when we encounter, in some iteration, an improvement smaller than some parameter $\beta>0$. Then, using a partial enumeration technique, we can eliminate this small cost again, see [8]. This is also the approach that [12] take.
[12] prove that, over a matroid constraint, the greedy algorithm would attain only approximation ratio $\frac{1}{2}$. Also, the approximation ratio of oblivious local search (Algorithm 1 in their paper)is $\frac{k-1}{2k-\ell-1}$ when $\ell$ sets are exchanged in each iteration. For $k=2$ and $\ell=1$, this also equals $\frac{1}{2}$, which is quite a lot smaller than $1-1/\operatorname{\text{e}}$. Thus, they switch to a non-oblivious local search algorithm that uses an auxiliary objective function $f$ for the iterative procedure. This auxiliary objective function adds temporary weight to elements covered multiple times. This makes a path of improvements more “robust”, and thereby prevents getting stuck in bad local optima. We denote by $\alpha_{j}$ the temporary weight associated to an element covered $j$ times, and write $h_{i}(W)$ for the amount of times that $i$ is covered by $W$. Then, the auxiliary objective function value of committee $W$ is
@@
Note that, with $\alpha_{0}=0$ and $\alpha_{j}=1$ for all $j>0$, we retain the original oblivious objective function, but if we set $\alpha_{j}>\alpha_{1}$ for $j>1$, we add additional weight to elements covered multiple times. As in [12], we define
$\alpha_{0}=0,\ \ \ \ \ \ \ \ \alpha_{1}=1-\frac{1}{\operatorname{\text{e}}},\ \ \ \ \ \ \ \ \alpha_{j+1}=(j+1)\alpha_{j}-j\alpha_{j-1}-\frac{1}{\operatorname{\text{e}}}.$ (5)
The sequence $(\alpha_{n})_{n\in\operatorname{\mathbb{N}}_{0}}$ thus gives zero points to a voter that is not covered by the committee at all, $1-\frac{1}{\operatorname{\text{e}}}\approx 0.632$ points to a voter that is covered by a single committee member, $2-\frac{3}{\operatorname{\text{e}}}\approx 0.896$ points to a voter that is covered twice, and so on. [12] prove this choice for the sequence $(\alpha_{n})_{n\in\operatorname{\mathbb{N}}_{0}}$ is optimal and that, for any $\gamma>0$, using this auxiliary objective function, the non-oblivious local search algorithm, with parameter $\beta$ (decreasing in $\gamma$), is $(1-1/\operatorname{\text{e}}-\gamma)$-approximate.
We have adjusted the non-oblivious local search algorithm of [12] to suit our setting better. We explain the changes below (in Section 4.2), as they are induced by the querying strategy. For now it is only important to know that the changes do not affect the approximation guarantees of the algorithm. See our adapted version, Algorithm 3, below, where, for $c\in W$ and $c^{\prime}\notin W$, we now write $\Delta(W,c^{\prime}):=f(W\cup\{c^{\prime}\})-f(W)$ and $\Delta(W,c^{\prime},c):=f((W\cup\{c^{\prime}\})\setminus\{c\})-f(W)$.
@@
Output: Committee $W$, consisting of $k$ candidates.
Choose $W\subset C$ such that $|W|=k$, and $c\in W$ and $c^{\prime}\notin W$ so that $(W\cup\{c^{\prime}\})\setminus\{c\}\in\mathcal{M}$;
@@
      until $\Delta(W,c^{\prime},c)\leq\beta$ and $f(W)>1/n$;
@@
[12, Corollary 6] prove that the local search algorithm with parameter $\beta=\frac{\gamma}{C_{0}k\log k(1-1/\operatorname{\text{e}}-\gamma)}$, is $(1-1/\operatorname{\text{e}}-\gamma)$-approximate for any $\gamma>0$ and some universal constant $C_{0}$ for which (by their Lemma 4) we know that $\alpha_{i}\leq C_{0}\log i$ for all $i\leq k$. Simplifying notation, we write $\beta=C_{1}\frac{\gamma}{k\log k}$ for any $\gamma>0$ and some universal constant $C_{1}=\frac{1}{C_{0}(1-1/\operatorname{\text{e}}-\gamma)}\leq\frac{\log i}{\alpha_{i}(1-1/\operatorname{\text{e}}-\gamma)}$ for any $i\leq k$.
The local search query algorithm
We turn to the incomplete information setting. Like before, we write $A:V\to\mathcal{P}(C)$ for the random variable, with unknown distribution $d$, that maps a voter $i$ to approval set $A(i)$. After querying any set $Q\subseteq C$ on all voters, we can compute, for any $S\subseteq Q$,
$p_{S}:=\mathbb{P}(A\cap Q=S)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(A(i)\cap Q=S),$ (6)
for the probability that a voter approves, within the set $Q$, of exactly the set $S$. Note that this is different than how $p_{S}$ was defined in Section 4.1, but both are meant to denote the relevant information that we attain from a query set $Q$. With this information, we can compute $f(Q)$. Since this time, we have, for $c^{\prime}\notin W$ and $c\in W$, $\Delta(W,c^{\prime}):=f(W\cup\{c^{\prime}\})-f(W)$ and $\Delta(W,c^{\prime},c):=f((W\cup\{c^{\prime}\})\setminus\{c\})-f(W)$, we can compute these values with complete information on a query set containing $W\cup\{c^{\prime}\}$. Indeed,
$\begin{split}f(W)&=\frac{1}{n}\sum_{i=1}^{n}\alpha_{h_{i}(S)}\\ &=\frac{1}{n}\sum_{i=1}^{n}\sum_{S\subseteq W}\mathds{1}(A(i)\cap W=S)\cdot\alpha_{|S|}\\ &=\sum_{S\subseteq W}\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(A(i)\cap W=S)\cdot\alpha_{|S|}\\ &=\sum_{S\subseteq W}p_{S}\cdot\alpha_{|S|}.\end{split}$ (7)
Sampling $\ell\leq n$ voters from $d$ independently and presenting them with query set $Q\subseteq C$, we can compute, for every subset $S\subseteq Q$,
$\hat{p}_{S}:=\frac{1}{\ell}\sum_{i=1}^{\ell}\mathds{1}(A(i)\cap Q=S),$ (8)
that estimates $p_{S}$. Consequently, we write $\hat{\Delta}(W,c^{\prime})=\hat{f}(W\cup\{c^{\prime}\})-\hat{f}(W)=\sum_{S\subseteq W\cup\{c^{\prime}\}}\hat{p}_{S}\cdot\alpha_{|S|}-\sum_{S\subseteq W}\hat{p}_{S}\cdot\alpha_{|S|}$ for an estimate of $\Delta(W,c^{\prime})$ based on the values $\hat{p}$) and $\hat{\Delta}(W,c^{\prime},c)$ as an estimate of $\Delta(W,c^{\prime},c)$ based on values $\hat{p}$ analogously. We present our querying algorithm in Algorithm 4.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C$ of $m$ candidates, $\beta>0$, query size $t$, with $m\geq t>k$, constants $\delta\in(0,1]$ and $\xi\geq 1$.
Output: Committee $W$, consisting of $k$ candidates.
@@
Choose $W\subset C$ such that $|W|=k$, and $c\in W$ and $c^{\prime}\notin W$ so that $(W\cup\{c^{\prime}\})\setminus\{c\}\in\mathcal{M}$;
@@
    Construct a set of query sets $\mathcal{Q}=\{Q_{i}\}_{i}$ with $Q_{i}\subseteq C$ and $|Q_{i}|=t$ for all $i$, and such that $W\subseteq\bigcap\mathcal{Q}$ and $\bigcup\mathcal{Q}=C$, and present each query set to $\ell$ voters.
@@
      until $\hat{\Delta}(W,c^{\prime},c)<\beta-\epsilon$ and $\hat{f}(W)>\alpha_{1}/n$;
ALGORITHM 4 Local Search-$\beta$-queries
As previously mentioned, in Algorithm 3 (hence also Algorithm 4) we implemented several changes with respect to Algorithm 2 of [12]. We explain them here. There are three changes. Firstly, we have left out the greedy initialization. Although it speeds up the convergence time-wise, it would cost many extra querying rounds. It is not possible to query only for the initialization and use the gathered information for the local search phase, since candidates not elected in the initialization, or only elected in its last round, will not have query responses from all voters whereas that is necessary if the candidate would become part of the tentative solution at some point during the local search. However, it might be possible to save the information gathered in the initialization in a smart way. For now, we leave out this phase. For similar reasons, we omit the partial enumeration step that would yield the clean $1-1/\operatorname{\text{e}}$ ratio. Secondly, we have adapted the stopping condition from $\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}\leq 1+\beta n$ (the factor $n$ is added because we use relative scores) to $f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta$. Before we can make any kind of statement about the performance of our algorithm, we need guarantees on the accuracy of our estimations. In order to use Hoeffding’s inequality to bound the probability that our estimates are far from the estimated value, we need our samples to be bounded as well. However, since $\alpha_{0}=0$, the expression $\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}$ is unbounded. Suppose now, that $f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta$, then
$f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta\implies\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}\leq 1+\frac{\beta}{f(W)},$ (9)
which is at most $1+\beta n$ iff $f(W)\geq 1/n$. This might not be true immediately, but will be true at some point as long as $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\geq 1$, so also when the instance is non-trivial. Thus, we also add the requirement that $\hat{f}(W)\geq\alpha_{1}/n$, and then our adapted stopping condition is a stronger requirement for any non-trivial instance. This makes our third adaptation. Theorem 4.10 gives the amount of queries required for Algorithm 4. As expected, this amount is higher than stated in Theorem 4.4 for the original problem.
Theorem 4.10.

### v2

@@ Introduction
ALGORITHM 2 greedy-incomplete
Theorem 2.
Let an instance be non-trivial, $\delta>0,\,\gamma\in(0,1)$, and $k<t\leq m$. Then, w.p. at least $1-\delta$, Algorithm 2 is $(1-1/e)\gamma$-approximate for Max Cover with query complexity $O\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)\in O_{\delta,\gamma,k}(m\log m)$.
Importantly, the query complexity does not depend on $n$. Asymptotically, $O(m\log m)$ is much smaller than querying the entire ballot of size $nm$. The constant overhead, which increases with $k$ and $\gamma$ and decreases with $\delta$, is quite small. For instance, if $m=1000,\,\gamma=0.85,\,\delta=0.05,\,k=8$ and $t=20$, Theorem 2 requires 432 920 queries. Starting with $n>432$ voters this will be smaller than $nm$.
Upper bound for a setting with matroid constraints
@@
We can imagine situations where we want to place restrictions on the set of possible committees to be elected. Perhaps we want to include extra diversity requirements, such as ‘the committee must contain at least $x$ candidates of category $Y$’. In practice, $x$ would be a natural number and $Y$ would be a category like a gender, race, etc. In online deliberation settings, these attributes would extend to users. For such purposes, we make a selection of which committees are “allowed” to be elected, and collect them in a set $\mathcal{M}$. This is fundamentally different from what we meant so far with diversity, which was defined in terms of the approval profile. It is not obvious that any algorithm adapts well to this restriction. Fortunately, when $\mathcal{M}$ is a matroid, we can mend this problem [11].
Definition 3 (matroid).
@@
Property 3) guarantees that all maximal independent sets have the same cardinality. We call such sets bases, and their common cardinality is called the rank of the matroid.
Example 4 (uniform matroid).
Consider $\mathcal{M}=\{W\subseteq C:|W|\leq k\}$. This is the uniform matroid of rank $k$: 1) $|\emptyset|=0\leq k$, 2) for any set $W\in\mathcal{M}$ with $|W|=j\leq k$, for any subset $W^{\prime}\subseteq W$, $|W^{\prime}|\leq|W|\leq k$, and 3) for any two feasible sets $W,\ W^{\prime}\in\mathcal{M}$ with $|W^{\prime}|<|W|$, we will have, for any $c\in W\setminus W^{\prime}$, that $|W^{\prime}\cup\{c\}|\leq k$.
Thus, taking for $\mathcal{M}$ the uniform matroid of rank $k$, we retrieve the original problem: the problem with matroid constraints is a generalization of the original problem.
Example 5 (matroid implementing quota).
For $X\cup Y=C,\ X\cap Y=\emptyset$, define a set of bases $M=\{W\subseteq C:|W|=k,\ \ |W\cap Y|=|W\cap X|=\frac{1}{2}k\}$. Then $\mathcal{M}=M\bigcup\ \{V:V\subseteq W,\ W\in M\}$ is a matroid of rank $k$.
In Example 5, only the bases comply with the quota, but this is not a problem. When an algorithm outputs $A\in\mathcal{M}$ with $|A|=\ell<k$, then by requirement 3) in Definition 3, using any basis element $B=\{b_{1},\ldots,b_{k}\}\in\mathcal{M}$, we can make a path $A,\ A\cup\{b_{1}\},\ A\cup\{b_{1}\}\cup\{b_{2}\},\ \ldots,A\cup\{b_{1}\}\cup\ldots\cup\{b_{k-\ell}\}=:A^{\prime}$ from AA to attain a basis element $A^{\prime}\in\mathcal{M}$ that contains AA (and has a score at least as high).
@@
As before, we start with the perfect information setting. A classical (or oblivious) local search algorithm for Max Cover starts with an arbitrary solution and, in each iteration, swaps one or multiple of the elected sets for unelected sets in order to improve the objective function value. It thus searches in the local neighborhood of the solution. It stops when no local improvement is possible. At worst, this occurs after all exponentially many options have been exhausted. With the use of approximate local search, this running time problem can be resolved at a small cost in the approximation ratio, see also [26]. In practice, this means we terminate the algorithm when we encounter an improvement smaller than some parameter $\beta>0$  [^3] Using a partial enumeration technique, we could even eliminate this small cost again, see [7]..
Over a matroid constraint, the greedy algorithm attains only approximation ratio $\frac{1}{2}<1-1/\operatorname{\text{e}}$. The approximation ratio of oblivious local search is $\frac{k-1}{2k-\ell-1}$ when $\ell$ sets are exchanged in each iteration [11]. For $k=2$ and $\ell=1$, this also equals $\frac{1}{2}$. Thus, we switch to a non-oblivious local search algorithm that uses an auxiliary objective function $f$ for the iterative procedure. The function adds temporary weight to elements covered multiple times, the idea being that elements covered multiple times, remain covered after the next exchange. Such a function can thus prevent getting stuck in bad local optima. We denote by $\alpha_{j}$ the temporary weight associated to an element covered $j$ times, and write $h_{i}(W)$ for the amount of times that $v_{i}$ is covered by $W$. The auxiliary objective function is defined
@@
Note that, with $\alpha_{0}=0$ and $\alpha_{j}=1$ for all $j>0$, we retain the original oblivious objective function, but if we set $\alpha_{j}>\alpha_{1}$ for $j>1$, we add additional weight to elements covered multiple times. We define
$\alpha_{0}=0,\ \ \ \ \alpha_{1}=1-\frac{1}{\operatorname{\text{e}}},\ \ \ \ \alpha_{j+1}=(j+1)\alpha_{j}-j\alpha_{j-1}-\frac{1}{\operatorname{\text{e}}}.$
This choice for the sequence $(\alpha_{n})_{n\in\operatorname{\mathbb{N}}_{0}}$ is optimal and, for any $\gamma\in(0,1)$, with this objective function, the non-oblivious local search algorithm, with parameter $\beta$ (decreasing in $\gamma$) is $(1-1/\operatorname{\text{e}}-\gamma)$-approximate and runs in polynomial time [11].
We adjust the non-oblivious local search algorithm of [11] to suit our setting better, see Algorithm 3, where for $c\in W$ and $c^{\prime}\notin W$, we now write $\Delta(W,c^{\prime},c):=f((W\cup\{c^{\prime}\})\setminus\{c\})-f(W)$. We elaborate on the adaptations in Appendix B, but most importantly, the changes do not affect the approximation guarantees of the algorithm: for $\beta=C_{1}\frac{\gamma}{k\log k}$, Algorithm 3 is $(1-1/\operatorname{\text{e}}-\gamma)$-approximate for any $\gamma\in(0,1)$ and some universal constant $C_{1}\leq\frac{\log i}{\alpha_{i}(1-1/\operatorname{\text{e}}-\gamma)}$ for any $i\leq k$ [11, Corollary 6].
@@
Output: Committee $W\subseteq C$, of size $k$.
Choose $W\subseteq C$ such that $|W|=k$, and $c\in W$ and $c^{\prime}\notin W$ so that $(W\cup\{c^{\prime}\})\setminus\{c\}\in\mathcal{M}$;
@@
   until $\Delta(W,c^{\prime},c)\leq\beta$;
@@
The non-oblivious local search query algorithm
We turn to the incomplete information setting. Recall that $A(i,j)$ equals 1 when voter $v_{i}$ approves of candidate $c_{j}$ and 0 otherwise, and that we write $A(i)$ for the set of candidates approved by voter $v_{i}$. Given a query set $Q\subseteq C$, for any $S\subseteq Q$, we redefine
$p_{S}:=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{A(i)\cap Q=S\}}(i)$ (5)
as the probability that a voter, of all the candidates in $Q$, approves of exactly the set $S$. This is different from how we defined $p_{S}$ for Algorithm 2: Algorithm 2 requires, for any voter $v_{i}$, knowing just whether they have a representative in set $S\subseteq Q$, whereas Algorithm 4 needs the number of representatives. With $p_{S}$ for $S\subseteq Q$, we can compute $f(Q)$:
$f(Q)$ $=\frac{1}{n}\sum_{i=1}^{n}\alpha_{h_{i}(Q)}=\frac{1}{n}\sum_{i=1}^{n}\sum_{S\subseteq W}\mathds{1}_{\{A(i)\cap Q=S\}}(i)\cdot\alpha_{|S|}$
$=\sum_{S\subseteq Q}\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{A(i)\cap Q=S\}}(i)\cdot\alpha_{|S|}=\sum_{S\subseteq Q}p_{S}\cdot\alpha_{|S|}.$
We can compute $\Delta(W,c^{\prime},c)$ with complete information on a query set containing $W\cup\{c^{\prime}\}$. When information is incomplete, we estimate: sampling $\ell\leq n$ voters uniformly at random and presenting them with query set $Q\subseteq C$, we can compute, for every subset $S\subseteq Q$,
$\hat{p}_{S}:=\frac{1}{\ell}\sum_{i=1}^{\ell}\mathds{1}_{\{A(i)\cap Q=S\}}(i),$ (6)
that estimates $p_{S}$. We write $\hat{\Delta}(W,c^{\prime},c)=\hat{f}((W\cup\{c^{\prime}\})\setminus\{c\})-\hat{f}(W)=\sum_{S\subseteq(W\cup\{c^{\prime}\})\setminus\{c\}}\hat{p}_{S}\cdot\alpha_{|S|}-\sum_{S\subseteq W}\hat{p}_{S}\cdot\alpha_{|S|}$ for an estimate of $\Delta(W,c^{\prime},c)$ based on the values $\hat{p}$. Algorithm 4 is our local search query algorithm. Theorem 6 gives the number of queries required for Algorithm 4. As expected, this is higher than stated in Theorem 2.
Input: Numbers $n,m,k\in\operatorname{\mathbb{N}}$ with $k\leq m$, set $V$ of $n$ voters, set $C$ of $m$ candidates, $\beta>0$, query size $t$, with $m\geq t>k$, constants $\delta>0$ and $\xi\geq 1$.
Output: Committee $W\subseteq C$, of size $k$.
@@
Choose $W\subseteq C$ such that $|W|=k$, and $c\in W$ and $c^{\prime}\notin W$ so that $(W\cup\{c^{\prime}\})\setminus\{c\}\in\mathcal{M}$;
@@
    Construct the smallest set of query sets $\mathcal{Q}=\{Q_{i}\}_{i}$ with $Q_{i}\subseteq C$ and $|Q_{i}|=t$ for all $i$, and such that $W\subseteq\bigcap\mathcal{Q}$ and $\bigcup\mathcal{Q}=C$, and present each query set to $\ell$ voters sampled uniformly at random.
@@
   until $\hat{\Delta}(W,c^{\prime},c)<\beta-\epsilon$;
ALGORITHM 4 ls-$\beta$-incomplete
Theorem 6.